OpenStudy (anonymous):

simplify: the little 3 over the square root of 32x^5y^9

OpenStudy (anonymous):

That's called a cube root. It's the same as raising the quantity to the power of 1/3. (A square root is to the power of 1/2) 32 = 2^5 32^(1/3) = 2^(5/3) (x^5)^(1/3) = x^(5/3) (y^9)^(1/3) = y^3 2^(5/3) * x^(5/3) * y^3 is the most simplified answer

OpenStudy (anonymous):

is that how it is written in radical notation?

OpenStudy (anonymous):

I'm not using radicals. I'm raising them to exponents to forgo any problems of communicating how to write that as a radical. Do you need it in radical notation?

OpenStudy (anonymous):

yes that is where i get confused.

OpenStudy (anonymous):

... is there a way to solve for radical notation?

OpenStudy (anonymous):

Sorry about not getting back to you. given the final answer of: 2^(5/3) * x^(5/3) * y^3 You can write the radical sign with the 3, denoting a cube root, with 32*x^5*y^9 within it, and that would be in radical form. However, you can further simplify it still while maintaining radical form. Such as: 2^(5/3) pulled out of the radical. Also, y^(9/3) is equal to y^3 So you have y^3 * 2^(5/3) * cube root(x^5). Or you can have y^3 * cuberoot(32*x^5). Choose whichever you think is most applicable in your case.

OpenStudy (anonymous):

THANKS! honestly your descriptions are helping me out!!

OpenStudy (anonymous):

You're welcome.